OpenStudy (anonymous):

factor completely. if polynomial is prime state this. -16x^2-32x-7

OpenStudy (anonymous):

-(4x+7)(4x+1)

OpenStudy (anonymous):

Do u want to show all all the work or are u fine???

OpenStudy (anonymous):

"prime" means "doesn't factor"

OpenStudy (anonymous):

if you don't mind could u please show work. that way i can try to do the next one on my own. thank you for helping me out.

OpenStudy (anonymous):

means write "PRIME" if you cannot factor

OpenStudy (anonymous):

Its k.I will give u an analogy u try with this for later probs too For any second order linear equation ax^2+bx+c=0; Just see that the middle one b must be factorised so that we can take common Now b should be factorised in such a way that adding the two factors must give the value of b and their product must be equal to product of a*c..... For this u just multiply a and c and then u just try to write the miltiple of the product and check whetther they add to give the value b.... Now here i will explain u with this prob see The given expression is -(16x^2+32x+7) Now 16 * 7= 4*4*7=28*4 Now u can see adding 28 and 4 gives 32. So the factors are 28 and 4. Now it can be written as -(16x^2+28x+4x+7) Implies -(4x(4x+7)+1(4x+7)) = -(4x+7)(4x+1) And thats the required factors

OpenStudy (anonymous):

thanks a lot i will remember that.

OpenStudy (anonymous):

No probs